OpenStudy (anonymous):

A wooden block slides down a 47 degree incline with acceleration 2.33 m/s2. Find the coefficient of kinetic friction.

OpenStudy (anonymous):

Firstly, we need Normal Contact force (N) and Force due to weight on inclined plane(Fw). Fw=mgsin$=m*10*sin(47)= 7.3m Newtons. n=mgcos$=m*10*cos(47)= 6.8m Newtons. Next stepfind the Frictional force(Ff), basing from newtons 2nd law of motion: Fw-Ff=ma 7.3m-Ff=m*2.33 Ff=4.97m Newtons Thirdly, using the formula Ff=£N, we will find £ or corfficient of friction Therefore 4.97m=£*6.8m, hence £= 0.73